# Total variation distance between SDEs with stable noise and Brownian motion with applications to Poisson PDEs

### v1

@@ 2.2. Integration by parts via Malliavin Calculus
$\|D_{v}J_{s,t}^{-1}\|_{\mathrm{op},\infty}\leq\|\sigma\|_{\mathrm{op}}\theta_{2}\|v\|_{H}\sqrt{t-\ell^{\epsilon}_{0}}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}},\quad s\in[\ell^{\epsilon}_{0},t].$
Then, it follows from the chain rule and (2.7) that
$\|D_{v}u_{i}\|^{2}_{H}$ $=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}}^{t}\left|\sigma^{-1}\left\{D_{v}J_{s,t}^{-1}\right\}e_{i}\right|^{2}\mathrm{d}s$
$\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}\|v\|_{H}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}.$
Since this holds for any $v\in H$, one can deduce that
$\|Du_{i}\|^{2}_{H\otimes H}\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}.$
@@
$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|$
$\quad\leq\left(\frac{\|\sigma^{-1}\|_{\mathrm{op}}}{\sqrt{t-\ell^{\epsilon}_{0}}}\,\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}+\|\sigma^{-1}\|_{\mathrm{op}}\|\sigma\|_{\mathrm{op}}\theta_{2}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big{|}^{2}\right]}$
$\quad\leq C\left(\frac{1}{\sqrt{t-\ell^{\epsilon}_{0}}}+\gamma^{\epsilon}_{t}\right)\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big{|}^{2}\right]}.$
@@
Assume (H2). For any $\alpha\in(0,2)$, $t\in(0,1]$, $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$ and $x\in{\mathds{R}^{d}}$,
[…]
@@
Letting $\epsilon\downarrow 0$ in Lemma 2.2 and using $\ell^{\epsilon}_{t}\downarrow\ell_{t}$ and [22, Lemma 2.2], we get for $t>0$,
$\left|\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\right|\leq C\left(\ell_{t}^{-1/2}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell}_{t})\right|^{2}\right]},\quad 1\leq i\leq d.$
Now we “unfreeze” the fixed path $\ell$ and use (2.2). From the Cauchy–Schwarz inequality we get for $t>0$,
$\left|P_{t}(\partial_{i}f)(x)\right|$ $=\left|\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)\right|$
$\leq\int_{\mathds{S}}\big{|}\mathds{E}^{\mathds{W}}\left[\partial_{i}f\left(X^{x;\ell}_{t}\right)\right]\big{|}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
$\leq C\mathrm{e}^{2\theta_{1}t}\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big{|}f(X^{x;\ell}_{t})\big{|}^{2}\right]}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
$\leq C\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)^{2}\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\big{|}f(X^{x;\ell}_{t})\big{|}^{2}\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)}$
$\leq C\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1}+t^{2}\right)\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\mathds{E}\left[\big{|}f(X^{x}_{t})\big{|}^{2}\right]}$
$=C\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}\left[S_{t}^{-1}\right]+t^{2}}\,\sqrt{P_{t}|f|^{2}(x)}$
$\leq C\mathrm{e}^{2\theta_{1}t}\left(\sqrt{\mathds{E}\left[S_{t}^{-1}\right]}+t\right)\sqrt{P_{t}|f|^{2}(x)}.$
@@
It follows from Lemma 2.3 that for all $t\in(0,1]$, $\alpha\in(0,2)$ and $x\in{\mathds{R}^{d}}$,
$\left\|\mathds{E}\left[\nabla^{2}f(X^{x}_{t})\right]\right\|_{\mathrm{HS}}^{2}$ $=\sum_{i,j=1}^{d}\left|\mathds{E}\left[\partial_{i}\partial_{j}f(X^{x}_{t})\right]\right|^{2}$
[…]
@@
$\sum_{j=1}^{d}P_{t}|\partial_{j}f|^{2}(x)=P_{t}|\nabla f|^{2}(x)\leq\|\nabla f\|_{\infty}^{2},$
@@ 3. Gradient estimates of $Q_{t}$ and comparison of the generators
    $\qquad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d^{3/2}+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d.$
@@
    $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\leq C_{1}d^{3/2}(2-\alpha)\left((t-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{e}^{-C_{2}k}.$
@@
=$=\left\{\int_{|z|<1}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
@@
=$=\left\{\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
@@
(3.4) $\begin{split}|P_{t-s}\mathsf{J}_{11}(x)|&=\left|\frac{1}{2}\left[\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right]\big{\langle}P_{t-s}\nabla^{2}f(x),\sigma\sigma^{\top}\big{\rangle}_{\mathrm{HS}}\right|\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{HS}}\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\\ &\leq C(2-\alpha)(t-s)^{-1/\alpha}\|\nabla f\|_{\infty}d\log(1+d).\end{split}$
Let us now estimate $P_{t-s}\mathsf{J}_{12}(x)$. Since $\|(\sigma z)(\sigma z)^{\top}\|_{\mathrm{HS}}=|\sigma z|^{2}$, it follows from Proposition 2.1 that for any $r\in[0,1]$ and $\alpha\in(1,2)$,
@@
$\quad\leq\left\|P_{t-s}\nabla^{2}\{f(\cdot+r\sigma z)-f(\cdot)\}(x)\right\|_{\mathrm{HS}}\left\|(\sigma z)(\sigma z)^{\top}\right\|_{\mathrm{HS}}$
$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\|\nabla\{f(\cdot+r\sigma z)-f(\cdot)\}\right\|_{\infty}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\{[2\|\nabla f\|_{\infty}]\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}r|\sigma z|]\right\}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\sqrt{d}\,\left\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\right\}|z|^{2}.$
@@
(3.5) $\begin{split}&|P_{t-s}\mathsf{J}_{12}(x)|\\ &\leq\int_{|z|<1}\int_{0}^{1}\left|\langle P_{t-s}\nabla^{2}f(x+r\sigma z)-P_{t-s}\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}\right|(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &\leq C(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\left(\int_{0}^{1}(1-r)\,\mathrm{d}r\right)\big{\{}\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big{\}}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &=\frac{C}{2}(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\big{\{}\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big{\}}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z.\end{split}$
@@
$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|<1}\left\{s^{-1/2}\wedge[s^{-1}|z|]\right\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$=C(t-s)^{-1/\alpha}\sqrt{d}\left(\int_{|z|\leq\sqrt{s}}s^{-1}|z|\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z+\int_{\sqrt{s}<|z|<1}s^{-1/2}\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z\right)$
$=C(t-s)^{-1/\alpha}\,s^{-1/2}\sqrt{d}\,A(d,\alpha)\omega_{d-1}\left(s^{-1/2}\int_{0}^{\sqrt{s}}r^{2-\alpha}\,\mathrm{d}r+\int_{\sqrt{s}}^{1}r^{1-\alpha}\,\mathrm{d}r\right)$
$=C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\sqrt{d}\,\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
$\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d^{3/2}.$
@@
$\quad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left[(2-\alpha)+\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\right]d^{3/2}+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d,$
@@
$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C_{1}(t-s)^{-1/\alpha}\sqrt{d}\int_{|z|\leq 1}\mathrm{e}^{-C_{2}k}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$=C_{1}(t-s)^{-1/\alpha}\sqrt{d}\,\mathrm{e}^{-C_{2}k}A(d,\alpha)\omega_{d-1}\int_{0}^{1}r^{2-\alpha}\,\mathrm{d}r$
$=C_{1}(t-s)^{-1/\alpha}\sqrt{d}\,\mathrm{e}^{-C_{2}k}\frac{A(d,\alpha)\omega_{d-1}}{3-\alpha}$
$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d^{3/2}.$
@@
$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d^{3/2}+\frac{C_{1}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k}d,$
@@
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq 2\wedge\left[\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}\right].$
@@
$\sup_{x\in{}^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}\sup_{x\in{}^{d}}\left|\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h(x)\,\mathrm{d}s\right|.$

### v2

@@ 2.2. Integration by parts via Malliavin Calculus
$\|D_{u}J_{s,t}^{-1}\|_{\mathrm{op},\infty}\leq\|\sigma\|_{\mathrm{op}}\theta_{2}\|u\|_{H}\sqrt{t-\ell^{\epsilon}_{0}}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}},\quad s\in[\ell^{\epsilon}_{0},t].$
Then, it follows from (2.7) that
$\|D_{u}u_{0}\|^{2}_{H}$ $=\frac{1}{(t-\ell^{\epsilon}_{0})^{2}}\int_{\ell^{\epsilon}_{0}}^{t}\left|\sigma^{-1}\left\{D_{u}J_{s,t}^{-1}\right\}v\right|^{2}\mathrm{d}s$
$\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}\|u\|_{H}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}|v|^{2}.$
Since this holds for any $u\in H$, one can deduce that
$\|Du_{0}\|^{2}_{H\otimes H}\leq\|\sigma^{-1}\|_{\mathrm{op}}^{2}\|\sigma\|_{\mathrm{op}}^{2}\theta_{2}^{2}(\gamma^{\epsilon}_{t})^{2}\mathrm{e}^{4\theta_{1}\gamma^{\epsilon}_{t}}|v|^{2}.$
@@
$\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\right]\right|$
$\quad\leq\left(\frac{\|\sigma^{-1}\|_{\mathrm{op}}|v|}{\sqrt{t-\ell^{\epsilon}_{0}}}\,\mathrm{e}^{\theta_{1}\gamma^{\epsilon}_{t}}+\|\sigma^{-1}\|_{\mathrm{op}}\|\sigma\|_{\mathrm{op}}\theta_{2}\gamma^{\epsilon}_{t}\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}|v|\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big|^{2}\right]}$
$\quad\leq C|v|\left(\frac{1}{\sqrt{t-\ell^{\epsilon}_{0}}}+\gamma^{\epsilon}_{t}\right)\mathrm{e}^{2\theta_{1}\gamma^{\epsilon}_{t}}\sqrt{\mathds{E}^{\mathds{W}}\left[\big|f\left(Z^{x;\ell^{\epsilon}}_{t}\right)\big|^{2}\right]}.$
@@
Assume (1.4) and (1.5). For any $\alpha\in(0,2)$, $t\in(0,1]$ and $f\in C^{1}_{b}({\mathds{R}^{d}},\real)$,
$\left\|P_{t}\nabla f\right\|_{\infty}\leq C\sqrt{\Gamma\left(1+\tfrac{2}{\alpha}\right)}\,2^{1/\alpha}t^{-1/\alpha}\|f\|_{\infty}.$
@@
Letting $\epsilon\downarrow 0$ in Lemma 2.2 and using $\ell^{\epsilon}_{t}\downarrow\ell_{t}$ and [24, Lemma 2.2], we get for $t>0$,
$\left|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell}_{t}\right)\right]\right|\leq C|v|\left(\ell_{t}^{-1/2}+t\right)\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}^{\mathds{W}}\left[\left|f(X^{x;\ell}_{t})\right|^{2}\right]}.$
Now we “unfreeze” the fixed path $\ell$ and use (2.2). From the Cauchy–Schwarz inequality we get for all $t>0$, $v\in\real^{d}$ and $x\in\real^{d}$,
$\left|P_{t}(\nabla_{v}f)(x)\right|$ $=\left|\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell}_{t}\right)\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)\right|$
$\leq\int_{\mathds{S}}\big|\mathds{E}^{\mathds{W}}\left[\nabla_{v}f\left(X^{x;\ell}_{t}\right)\right]\big|\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
$\leq C|v|\mathrm{e}^{2\theta_{1}t}\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)\sqrt{\mathds{E}^{\mathds{W}}\left[\big|f(X^{x;\ell}_{t})\big|^{2}\right]}\,\mu_{\mathds{S}}(\mathrm{d}\ell)$
$\leq C|v|\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1/2}+t\right)^{2}\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\int_{\mathds{S}}\mathds{E}^{\mathds{W}}\left[\big|f(X^{x;\ell}_{t})\big|^{2}\right]\,\mu_{\mathds{S}}(\mathrm{d}\ell)}$
$\leq C|v|\mathrm{e}^{2\theta_{1}t}\sqrt{\int_{\mathds{S}}\left(\ell_{t}^{-1}+t^{2}\right)\,\mu_{\mathds{S}}(\mathrm{d}\ell)}\,\sqrt{\mathds{E}\left[\big|f(X^{x}_{t})\big|^{2}\right]}$
$=C|v|\mathrm{e}^{2\theta_{1}t}\sqrt{\mathds{E}\left[S_{t}^{-1}\right]+t^{2}}\,\sqrt{P_{t}|f|^{2}(x)}$
$\leq C|v|\mathrm{e}^{2\theta_{1}t}\left(\sqrt{\mathds{E}\left[S_{t}^{-1}\right]}+t\right)\|f\|_{\infty}.$
@@
It follows from Lemma 2.3 that for all $t\in(0,1]$ and $\alpha\in(0,2)$,
$\|P_{t}\nabla^{2}f\|_{\mathrm{op},\infty}$ $=\sup_{x\in\real^{d}}\sup_{v_{1},v_{2}\in\real^{d},|v_{1}|=|v_{2}|=1}\left|P_{t}\nabla_{v_{2}}\nabla_{v_{1}}f(x)\right|$
$\leq C\sqrt{\Gamma\left(1+\tfrac{2}{\alpha}\right)}\,2^{1/\alpha}t^{-1/\alpha}\sup_{v_{1}\in\real^{d},|v_{1}|=1}\|\nabla_{v_{1}}f\|_{\infty}.$
@@
$\sup_{v\in\real^{d},|v|=1}\|\nabla_{v}f\|_{\infty}=\sup_{x\in\real^{d}}\sup_{v\in\real^{d},|v|=1}\left|\langle\nabla f(x),v\rangle\right|=\|\nabla f\|_{\infty},$
@@ 3. Gradient estimates of $Q_{t}$ and comparison of the generators
In the following lemma, we will frequently need the constants
$A(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{2-\alpha}\pi^{d/2}\Gamma(1-\frac{\alpha}{2})},\quad\omega_{d-1}:=\frac{2\pi^{d/2}}{\Gamma\left(\frac{d}{2}\right)},$
which can be estimated as follows (see [5, Lemma 3.3]):
@@
    $\qquad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(3-\alpha-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d\log(1+d)+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d.$
@@
    $\sup_{h\in\mathscr{C}_{b}(1)}\left\|P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s+k}h\right\|_{\infty}\leq C_{1}(2-\alpha)\left((t-s)^{-1/\alpha}+\frac{1}{\alpha-1}\right)\mathrm{e}^{-C_{2}k}d\log(1+d).$
@@
$=\left\{\int_{|z|<1}\left[f(x+\sigma z)-f(x)-\langle\nabla f(x),\sigma z\rangle\right]\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
@@
$=\left\{\int_{|z|<1}\int_{0}^{1}\langle\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}(1-r)\,\mathrm{d}r\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z-\frac{1}{2}\,\langle\nabla^{2}f(x),\sigma\sigma^{\top}\rangle_{\mathrm{HS}}\right\}$
@@
(3.5) $\begin{split}|P_{t-s}\mathsf{J}_{11}(x)|&=\left|\frac{1}{2}\left[\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right]\big\langle P_{t-s}\nabla^{2}f(x),\sigma\sigma^{\top}\big\rangle_{\mathrm{HS}}\right|\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{HS}}\|\sigma\sigma^{\top}\|_{\mathrm{HS}}\\ &\leq\frac{1}{2}\left|\frac{A(d,\alpha)\omega_{d-1}}{d(2-\alpha)}-1\right|d\|P_{t-s}\nabla^{2}f(x)\|_{\mathrm{op},\infty}\|\sigma\|_{\mathrm{op}}^{2}\\ &\leq C(2-\alpha)(t-s)^{-1/\alpha}\|\nabla f\|_{\infty}d\log(1+d).\end{split}$
Let us now estimate $P_{t-s}\mathsf{J}_{12}(x)$. It follows from Proposition 2.1 that for any $r\in[0,1]$ and $\alpha\in(1,2)$,
@@
$\quad\leq\left\|P_{t-s}\nabla^{2}\{f(\cdot+r\sigma z)-f(\cdot)\}(x)\right\|_{\mathrm{op}}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\left\|\nabla\{f(\cdot+r\sigma z)-f(\cdot)\}\right\|_{\infty}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\left\{[2\|\nabla f\|_{\infty}]\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}r|\sigma z|]\right\}|\sigma z|^{2}$
$\quad\leq C(t-s)^{-1/\alpha}\left\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\right\}|z|^{2}.$
@@
(3.6) $\begin{split}&|P_{t-s}\mathsf{J}_{12}(x)|\\ &\leq\int_{|z|<1}\int_{0}^{1}\left|\langle P_{t-s}\nabla^{2}f(x+r\sigma z)-P_{t-s}\nabla^{2}f(x),(\sigma z)(\sigma z)^{\top}\rangle_{\mathrm{HS}}\right|(1-r)\,\mathrm{d}r\,\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &\leq C(t-s)^{-1/\alpha}\int_{|z|<1}\left(\int_{0}^{1}(1-r)\,\mathrm{d}r\right)\big\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big\}|z|^{2}\frac{A(d,\alpha)}{|z|^{d+\alpha}}\,\mathrm{d}z\\ &=\frac{C}{2}(t-s)^{-1/\alpha}\int_{|z|<1}\big\{\|\nabla f\|_{\infty}\wedge[\|\nabla^{2}f\|_{\mathrm{op},\infty}|z|]\big\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z.\end{split}$
@@
$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C(t-s)^{-1/\alpha}\int_{|z|<1}\left\{s^{-1/2}\wedge[s^{-1}|z|]\right\}\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$=C(t-s)^{-1/\alpha}\left(\int_{|z|\leq\sqrt{s}}s^{-1}|z|\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z+\int_{\sqrt{s}<|z|<1}s^{-1/2}\,\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z\right)$
$=C(t-s)^{-1/\alpha}\,s^{-1/2}A(d,\alpha)\omega_{d-1}\left(s^{-1/2}\int_{0}^{\sqrt{s}}r^{2-\alpha}\,\mathrm{d}r+\int_{\sqrt{s}}^{1}r^{1-\alpha}\,\mathrm{d}r\right)$
$=C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\frac{A(d,\alpha)\omega_{d-1}}{2-\alpha}$
$\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)d.$
@@
$\quad\leq C(t-s)^{-1/\alpha}\,s^{-1/2}\left[(2-\alpha)+\left(1-\frac{1}{3-\alpha}\,s^{1-\frac{\alpha}{2}}\right)\right]d\log(1+d)+\frac{C}{\alpha-1}(2-\alpha)s^{-1/2}d,$
@@
$|P_{t-s}\mathsf{J}_{12}(x)|$ $\leq C_{1}(t-s)^{-1/\alpha}\int_{|z|\leq 1}\mathrm{e}^{-C_{2}k}|z|\frac{A(d,\alpha)}{|z|^{d+\alpha-2}}\,\mathrm{d}z$
$=C_{1}(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}A(d,\alpha)\omega_{d-1}\int_{0}^{1}r^{2-\alpha}\,\mathrm{d}r$
$=C_{1}(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}\frac{A(d,\alpha)\omega_{d-1}}{3-\alpha}$
$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d.$
@@
$\leq C_{1}(2-\alpha)(t-s)^{-1/\alpha}\mathrm{e}^{-C_{2}k}d\log(1+d)+\frac{C_{1}}{\alpha-1}(2-\alpha)\mathrm{e}^{-C_{2}k}d,$
@@
$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}\leq\frac{Cd^{3/2}}{\alpha-1}(2-\alpha)\cdot\left\{\left(t^{-\frac{1}{\alpha}+\frac{1}{2}}\log\tfrac{1}{t}\right)\vee 1\right\}.$
@@
$\sup_{x\in\real^{d}}\|\mathrm{law}(X_{t}^{x})-\mathrm{law}(Y_{t}^{x})\|_{\mathrm{TV}}=\sup_{h\in\mathscr{C}_{b}(1)}\sup_{x\in\real^{d}}\left|\int_{0}^{t}P_{t-s}(\mathscr{A}^{P}-\mathscr{A}^{Q})Q_{s}h(x)\,\mathrm{d}s\right|.$